Find the amplitude, wavelength, period, frequency, speed and displacement after 2.0

seconds of wave equation y(x,t) = 2 sin(3x − 4t). ( where x and y is in meter)


1
Expert's answer
2021-06-03T09:20:02-0400
"A=2\ m\\f=\frac{4}{2\pi}=0.64\ Hz\\T=1.57\ s\\v=\frac{4}{3}=1.33\frac{m}{s}"